OpenStudy (anonymous):

Tony bought a desktop computer and a laptop computer. Before finance charges, the laptop cost $400 less than the desktop. He paid for the computers using two different financing plans. For the desktop the interest rate was 8.5% per year, and for the laptop it was 5% per year. The total finance charges for one year were $412 . How much did each computer cost before finance charges?

OpenStudy (dtan5457):

Might have to give me 10 minutes for this one..

OpenStudy (dtan5457):

Before finance charges, the laptop cost $400 less than the desktop. desktop: x laptop: x-400 interest on desktop: x(0.085) interest on laptop: x-400(0.05) total finance charges were $412 now you need to add the interests 0.085x+0.05(x-400)=412 multiply everything by 100 85x+50x-50(400)=412000 90x-20000=412000 add the 20000 90x=432000 x=4800<<cost of desktop 4800-400=4400<<cost of laptop

OpenStudy (dtan5457):

Good?

OpenStudy (anonymous):

thanks